-2x(x+3)-(x+1)(x-2)=
-2x^2-5x-x+1
--2x^2+6x+1
is this right??

Answer by rothauserc(4718)

-2x(x+3)-(x+1)(x-2)=
-2x^2 -6x - (x^2-x-2) =
-x^2 -6x -x^2 +x +2
-5x+2
order of operations:
do everything in parenthesis first
evaluate by doing all multiplication and division from left to right, then addition and subtraction from left to right